# Integrating Dynamical Systems Modeling with Spatiotemporal scRNA-seq Data Analysis

### v1

@@ 3.2.1 Discrete Temporal Dynamics Modeling
Unbalanced Optimal Transport
Static optimal transport inherently conserves mass. However, during cellular development and differentiation, processes such as cell proliferation and apoptosis result in mass non-conservation. Consequently, it is essential to consider unnormalized distributions that account for cell growth and death. Additionally, the marginals must be adjusted to incorporate these factors effectively. So the unbalanced optimal transport can be defined as follows:
$\min_{\boldsymbol{\pi}\in\mathbb{R}_{+}^{N\times M}}\langle\boldsymbol{\pi},\mathbf{c}\rangle+\tau_{1}\operatorname{KL}(\boldsymbol{\pi}\mathbf{1}_{M}||\nu_{0})+\tau_{2}\operatorname{KL}(\boldsymbol{\pi}^{T}\mathbf{1}_{N}||\nu_{1}),$ (3.8)
where $\tau_{1}$ and $\tau_{2}$ are hyperparameters that control the degree of penalization. When $\tau_{1}=\tau_{2}\rightarrow+\infty$, then one can recover the original optimal transport. To further adjust the marginal distributions accounting for growth and death (Klein et al., 2025; Schiebinger et al., 2019), for the left marginal distribution $\nu_{0}$ we set
$(\nu_{0})_{i}=\frac{g\left(\mathbf{x}_{i}\right)^{t_{2}-t_{1}}}{\sum_{j=1}^{N}g\left(\mathbf{x}_{j}\right)^{t_{2}-t_{1}}},\quad\forall i\in\{1,\ldots,N\}.$ (3.9)
where $g$ is the growth/death function and can be estimated through the gene sets. For the right marginal distribution $\nu_{1}$, set it as the uniform distribution, i.e., $(\nu_{1})_{j}=1/M,\quad\forall j\in\{1,\ldots,M\}.$
By the obtained coupling matrix $\boldsymbol{\pi}\in\mathbb{R}_{+}^{N\times M}$, one can perform biological downstream analysis such as compute ancestors or descends of a cell state and impute gene expressions (Schiebinger et al., 2019; Klein et al., 2025). Naturally, a Markov Chain model can be formulated to quantify the transition probability among cells across time points based on the optimal transport plan (Klein et al., 2025; Weiler et al., 2024). By weighting this random walk with those induced by other quantities such as gene expression similairity (e.g. DiffuionMap), pseudotime (e.g. Palantir) or RNA Velocity, the CellRank analysis could also be conducted to dissect the underlying structure of the transitional dynamics (Weiler et al., 2024).
@@ 3.2.2 Continuous Temporal Dynamics Modeling
Although static optimal transport provides a robust framework for coupling distributions at different time points, there is a substantial interest in capturing continuous cellular dynamics over time and fitting mechanistic models that transform the source distribution into the target distribution. This interest has driven the development of various dynamical optimal transport (OT) methods and flow-based generative models. Prominent approaches include those based on the Benamou-Brenier formulation (Benamou and Brenier, 2000), such as TrajectoryNet (Tong et al., 2020), MIOFlow (Huguet et al., 2022), and other related methodologies (Ruthotto et al., 2020; Liu et al., 2021; Cheng et al., 2024a; Wan et al., 2023; Pooladian et al., 2024; Zhang et al., 2024a). Additionally, unbalanced dynamic OT methods (Peng et al., 2024; Sha et al., 2024; Tong et al., 2023; Eyring et al., 2024), Gromov-Wasserstein OT approaches (Klein et al., 2023), continuous normalizing flows (CNF), and conditional flow matching techniques (CFM) (Lipman et al., 2023; Tong et al., 2024a; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023b; Jiao et al., 2024b; Gao et al., 2024; Liu et al., 2022b; Wu et al., 2023; Jin et al., 2024; Chow et al., 2020; Cheng et al., 2024b) have also been proposed. Despite these advancements, many of these methods do not fully account for stochastic dynamical effects, particularly the intrinsic noise inherent in gene expression and cell differentiation (Zhou et al., 2021a, b), which are prevalent in single-cell biological processes (Elowitz et al., 2002).
@@
In the realm of stochastic dynamics, the Schrödinger bridge (SB) problem seeks to identify the most probable stochastic transition path between two arbitrary distributions relative to a reference stochastic process (Léonard, 2014). Variants of the SB problem have been applied across various domains, including single-cell RNA sequencing (scRNA-seq) analysis and generative modeling. These approaches encompass static methods (Lavenant et al., 2024; Pariset et al., 2023; Ventre et al., 2023; Chizat et al., 2022; Shi et al., 2024; De Bortoli et al., 2021; Pooladian and Niles-Weed, 2024; Liu et al., 2022a; Gu et al., 2024a), dynamic methods (Koshizuka and Sato, 2023; Neklyudov et al., 2023, 2024; Zhang et al., 2024b; Bunne et al., 2023a; Chen et al., 2022b; Albergo et al., 2023; Wang et al., 2021a; Jiao et al., 2024a; Zhou et al., 2024a; Liu et al., 2023a; Zhou et al., 2024b; Zhu et al., 2024b; Maddu et al., 2024; Yeo et al., 2021; Jiang and Wan, 2024), and flow-matching techniques (Tong et al., 2024b). However, these methods often fail to address unnormalized distributions resulting from cell growth and death. To address this problem, some methods have been developed to account for the unbalanced stochastic dynamics, for example, those based on branching SDE theory (e.g, gWOT) (Chizat et al., 2022; Zhang et al., 2021; Ventre et al., 2023; Lavenant et al., 2024) and those based on the Feynman-Kac formula with forward-backward SDE theory (Pariset et al., 2023). Among those, most methods often require prior knowledge of these processes, such as growth or death rates (Pariset et al., 2023; Lavenant et al., 2024; Chizat et al., 2022; Zhang et al., 2021) or depend on additional information like cell lineage data (Ventre et al., 2023).
@@
Recently, regularized unbalanced optimal transport (RUOT), also known as unbalanced Schrödinger bridge (Chen et al., 2022c), has emerged as a promising approach for modeling stochastic unbalanced continuous dynamics (Baradat and Lavenant, 2021; Chen et al., 2022c; Buze and Duong, 2023; Janati et al., 2020). RUOT can be viewed as an unbalanced relaxation of the dynamic Schrödinger bridge formulation. For instance, (Baradat and Lavenant, 2021) elucidates the connection between certain RUOT formulations and branching Schrödinger bridges. Meanwhile, a new deep learning framework (DeepRUOT) (Zhang et al., 2025), has been developed to learn general RUOT and infer continuous unbalanced stochastic dynamics from sample data based on derived Fisher regularization forms without requiring prior knowledge.
Figure 3: Dynamic Modeling of Temporally-resolved Single-cell Transcriptomics
@@
The primary objective now transforms to determine the dynamics described by Eq. (2.2) and Eq. (2.3) from observed data, given unnormalized distributions at $T$ discrete time points where $\mathbf{x}_{i}\in\mathbb{R}^{d}\sim{\nu}_{i}$ for each fixed time point $i\in\{0,\ldots,T-1\}$. Note that solely satisfying Eq. (2.3) does not admit a unique solution. Consequently, we need to ensure that the inferred dynamics also adhere to certain energy minimization principles. Building upon Eq. (2.3), the problem can be categorized into four distinct scenarios: (1) ${g}(\mathbf{x},t)=0\text{ and }\boldsymbol{\sigma}(\mathbf{x},t)=0,$ (2) ${g}(\mathbf{x},t)=0\text{ and }\boldsymbol{\sigma}(\mathbf{x},t)\neq 0,$ (3) ${g}(\mathbf{x},t)\neq 0\text{ and }\boldsymbol{\sigma}(\mathbf{x},t)=0,$ (4) ${g}(\mathbf{x},t)\neq 0\text{ and }\boldsymbol{\sigma}(\mathbf{x},t)\neq 0.$ Each of these cases is examined to systematically address the learning of the underlying dynamics.
Dynamical Optimal Transport ($g=0,\sigma=0$)
@@
In this case, it means the dynamics do not account for unblancedness and stochastic, i,e, the cellular dynamics is governed by $\mathrm{d}{\mathbf{x}}_{t}=\mathbf{b}({\mathbf{x}}_{t},t)\mathrm{d}t$. Then we can use the dynamical optimal transport to model these dynamics also known as Benamou-Brenier formulation (Benamou and Brenier, 2000), which can be stated as follows.
@@
$\begin{gathered}\dfrac{1}{2}\mathcal{W}_{2}^{2}\left(\nu_{0},\nu_{1}\right)=\inf_{\left({p}(\mathbf{x},t),\mathbf{b}(\mathbf{x},t)\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\|\mathbf{b}(\mathbf{x},t)\|_{2}^{2}{p}(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t,\\ \mbox{s.t. }\partial_{t}p+\nabla\cdot\left(\mathbf{b}(\mathbf{x},t)p\right)=0,\ p|_{t=0}=\nu_{0},p|_{t=1}=\nu_{1}.\end{gathered}$ (3.10)
@@
The inclusion of the factor $\frac{1}{2}$ on the left-hand side ensures that the Wasserstein distance has a more physically meaningful interpretation; for example, it represents the total action required to transport one distribution to another. In this formulation, probability distributions are connected through a deterministic transport equation. It has been demonstrated that this dynamic formulation is equivalent to static optimal transport problem (Problem (3.2.1)) when employing the cost function $c(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$
Nerual ODE Solver
@@
Numerous methodologies have been proposed to solve dynamical optimal transport or its variants numerically. The basic approach involves employing a neural network, denoted as $\mathbf{b}_{\theta}(\mathbf{x},t)$ , to parameterize $\mathbf{b}(\mathbf{x},t)$ , and subsequently utilizing the ordinary differential equations (ODEs) that govern particle trajectories. From Problem (3.10), it is evident that the optimization process must address two distinct loss components: the first pertains to the computation of an energy-related loss, while the second concerns the reconstruction error ($p(\mathbf{x},1)=\nu_{1}$).
@@
Regarding energy loss, the high-dimensional nature of the integral presents significant challenges due to the curse of dimensionality. To mitigate this issue, the integral is approximated using Monte Carlo integration and continuous normalizing flows (CNF) techniques (Tong et al., 2020). The strategy involves performing integration along the particle trajectories dictated by the ODE, i.e.,
$\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\|\mathbf{b}(\mathbf{x},t)\|_{2}^{2}{p}(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t=\mathbb{E}_{\mathbf{x}_{0}\sim\nu_{0}}\int_{0}^{1}\frac{1}{2}\|\mathbf{b}(\mathbf{x}(t),t)\|_{2}^{2}\mathrm{d}t,$
@@
where $\mathbf{x}(t)$ satisfies the ODE $\frac{\mathrm{d}\mathbf{x}}{\mathrm{d}t}=\mathbf{b}(\mathbf{x},t),\mathbf{x}_{0}\sim\nu_{0}$. For the distribution reconstruction loss, the authors incorporate an additional penalizing constraint. By integrating these two loss components, there exists a sufficiently large $\lambda\geq 0$ such that (Tong et al., 2020; Huguet et al., 2022)
@@
$\dfrac{1}{2}\mathcal{W}_{2}^{2}\left(\nu_{0},\nu_{1}\right)=\inf_{({p}(\mathbf{x},t),\mathbf{b}(\mathbf{x},t))}\mathbb{E}_{\mathbf{x}_{0}\sim\nu_{0}}\int_{0}^{1}\frac{1}{2}\|\mathbf{b}(\mathbf{x}(t),t)\|_{2}^{2}\mathrm{d}t+\lambda\mathcal{D}(p(\mathbf{x},1),\nu_{1}).$
@@
Based on this formulation, TrajectoryNet (Tong et al., 2020) computes both the energy and the reconstruction errors by using neural ODE (Chen et al., 2018a) to parametrize the velocity $\mathbf{b}(\mathbf{x},t)$.
@@
Recently, conditional flow matching (CFM) presents another efficient dynamical OT solver especially in high dimensionality case (Lipman et al., 2023; Tong et al., 2024a; Liu et al., 2023b; Albergo et al., 2023). Assume that the probability path ${p}(\mathbf{x},t)$ and the corresponding vector field $\mathbf{b}(\mathbf{x},t)$ generating it are known, and that ${p}(\mathbf{x},t)$ can be efficiently sampled. Under these conditions, a neural network $\mathbf{b}_{\theta}(\mathbf{x},t)$ can be trained to approximate $\mathbf{b}(\mathbf{x},t)$ by minimizing the flow matching (FM) objective:
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1),\mathbf{x}\sim{p}(\mathbf{x},t)}\|\mathbf{b}_{\theta}(\mathbf{x},t)-\mathbf{b}(\mathbf{x},t)\|_{2}^{2}.$
@@
However, this objective is computationally intractable when dealing with general source and target distributions. Consider the specific case of Gaussian marginal densities, defined as ${p}(\mathbf{x},t)=\mathcal{N}(\mathbf{x}|\mathbf{\mu}(t),\sigma(t)^{2}\mathbf{I})$. The corresponding unique vector field that generate this density from $\mathcal{N}(\mathbf{x}|\mathbf{\mu}(0),\sigma(0)^{2}\mathbf{I})$ is $\mathbf{b}(\mathbf{x},t)=\mathbf{\mu}^{\prime}(t)+\frac{\sigma(t)}{\sigma^{\prime}(t)}(\mathbf{x}-\mathbf{\mu}(t)),$ where $\mu^{\prime}(t)$ and $\sigma^{\prime}(t)$ means the time derivative (Tong et al., 2024a; Lipman et al., 2023). Now assume the marginal probability trajectory ${p}(\mathbf{x},t)$ is a mixture of conditional probability paths $p(\mathbf{x},t|\mathbf{z})$. Specifically, this can be expressed as:
${p}(\mathbf{x},t)=\int p(\mathbf{x},t|\mathbf{z})q(\mathbf{z})\mathrm{d}\mathbf{z}.$
@@
If the $p(\mathbf{x},t|\mathbf{z})$ is generated by the vector field $\mathbf{b}(\mathbf{x},t|\mathbf{z})$ from $p(\mathbf{x},0|\mathbf{z})$, then ${p}(\mathbf{x},t)$ can be generated by $\mathbf{b}(\mathbf{x},t)$ defined as follows:
$\mathbf{b}(\mathbf{x},t):=\mathbb{E}_{q}(\mathbf{z})\frac{\mathbf{b}(\mathbf{x},t|\mathbf{z})p(\mathbf{x},0|\mathbf{z})}{{p}(\mathbf{x},t)}.$
@@
This is also intractable since ${p}(\mathbf{x},t)$ is difficult to compute. The key is to introduce the conditional flow matching objective:
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1),q(\mathbf{z}),p(\mathbf{x},t|\mathbf{z})}\|\mathbf{b}_{\theta}(\mathbf{x},t)-\mathbf{b}(\mathbf{x},t|\mathbf{z})\|_{2}^{2}.$
@@
One can prove that $\nabla_{\theta}\mathcal{L}_{\text{CFM}}=\nabla_{\theta}\mathcal{L}_{\text{FM}}$, so training with CFM is equivalent with FM. The CFM objective is very useful when the $\mathbf{b}(\mathbf{x},t)$ is intractable but the conditional $\mathbf{b}(\mathbf{x},t|\mathbf{z})$ is tractable. So to approximate the dynamical optimal transport (3.10) is to use CFM. Assume $q(\mathbf{z})=q(\mathbf{z}_{0},\mathbf{z}_{1})$, and set $q(\mathbf{z})$ to be the Wasserstein optimal transport map $\boldsymbol{\pi}$ between the source distribution $\nu_{0}$ and the target distribution $\nu_{1}$, i.e., $q(\mathbf{z})=\boldsymbol{\pi}(\mathbf{z}_{0},\mathbf{z}_{1})$, where $\mathbf{z}_{0}\sim\nu_{0}$, $\mathbf{z}_{1}\sim\nu_{1}$. Then one can construct the Gaussian flow between $\mathbf{z}_{0}$ and $\mathbf{z}_{1}$ with standard deviation $\sigma$,
$p(\mathbf{x},t|\mathbf{z})=\mathcal{N}(\mathbf{x}|t\mathbf{z}_{1}+(1-t)\mathbf{z}_{0}|\sigma^{2}),\mathbf{b}(\mathbf{x},t|\mathbf{z})=\mathbf{z}_{1}-\mathbf{z}_{0}.$

### v2

@@ 3.2.1 Discrete Temporal Dynamics Modeling
Unbalanced Optimal Transport Static optimal transport inherently conserves mass. However, during cellular development and differentiation, processes such as cell proliferation and apoptosis result in mass non-conservation. Consequently, it is essential to consider unnormalized distributions that account for cell growth and death. Additionally, the marginals must be adjusted to incorporate these factors effectively. So, the unbalanced optimal transport can be defined as follows:
$\min_{\boldsymbol{\pi}\in\mathbb{R}_{+}^{N\times M}}\langle\boldsymbol{\pi},\boldsymbol{c}\rangle+\tau_{1}\operatorname{KL}(\boldsymbol{\pi}\boldsymbol{1}_{M}||\nu_{0})+\tau_{2}\operatorname{KL}(\boldsymbol{\pi}^{T}\boldsymbol{1}_{N}||\nu_{1}),$ (9)
where $\tau_{1}$ and $\tau_{2}$ are hyperparameters that control the degree of penalization. When $\tau_{1}=\tau_{2}\rightarrow+\infty$, then one can recover the original optimal transport. To further adjust the marginal distributions accounting for growth and death Klein et al. (2025); Schiebinger et al. (2019), for the left marginal distribution $\nu_{0}$ we set
$(\nu_{0})_{i}=\frac{g\left(\boldsymbol{x}_{i}\right)^{t_{2}-t_{1}}}{\sum_{j=1}^{N}g\left(\boldsymbol{x}_{j}\right)^{t_{2}-t_{1}}},\quad\forall i\in\{1,\ldots,N\}.$ (10)
where $g$ is the growth/death function and can be estimated through the gene sets. For the right marginal distribution $\nu_{1}$, set it as the uniform distribution, i.e., $(\nu_{1})_{j}=1/M,\linebreak\quad\forall j\in\{1,\ldots,M\}.$
By the obtained coupling matrix $\boldsymbol{\pi}\in\mathbb{R}_{+}^{N\times M}$, one can perform biological downstream analysis such as computing ancestors or descends of a cell state and imputing gene expressions Schiebinger et al. (2019); Klein et al. (2025). Naturally, a Markov chain model can be formulated to quantify the transition probability among cells across time points based on the optimal transport plan Klein et al. (2025); Weiler et al. (2024). By weighting this random walk with those induced by other quantities such as gene expression similairity (e.g., DiffuionMap), pseudotime (e.g., Palantir), or RNA velocity, the CellRank analysis could also be conducted to dissect the underlying structure of the transitional dynamics Weiler et al. (2024).
@@ 3.2.2 Continuous Temporal Dynamics Modeling
Although static optimal transport provides a robust framework for coupling distributions at different time points, there is a substantial interest in capturing continuous cellular dynamics over time and fitting mechanistic models that transform the source distribution into the target distribution. This interest has driven the development of various dynamical optimal transport (OT) methods and flow-based generative models. Prominent approaches include those based on the Benamou–Brenier formulation Benamou and Brenier (2000), such as TrajectoryNet Tong et al. (2020), MIOFlow Huguet et al. (2022), and other related methodologies Ruthotto et al. (2020); Liu et al. (2021); Cheng et al. (2024); Wan et al. (2023); Pooladian et al. (2024); Zhang et al. (2024). Additionally, unbalanced dynamic OT methods Peng et al. (2024); Sha et al. (2024); Tong et al. (2023); Eyring et al. (2024), Gromov–Wasserstein OT approaches Klein et al. (2023), continuous normalizing flows (CNF), and conditional flow matching techniques (CFM) Lipman et al. (2023); Tong et al. (2024); Albergo and Vanden-Eijnden (2023); Liu et al. (2023); Jiao et al. (2024); Gao et al. (2024); Liu et al. (2022); Wu et al. (2023); Jin et al. (2024); Chow et al. (2020); Cheng et al. (2024) have also been proposed. Despite these advancements, many of these methods do not fully account for stochastic dynamical effects, particularly the intrinsic noise inherent in gene expression and cell differentiation Zhou et al. (2021, 2021), which are prevalent in single-cell biological processes Elowitz et al. (2002).
@@
In the realm of stochastic dynamics, the Schrödinger bridge (SB) problem seeks to identify the most probable stochastic transition path between two arbitrary distributions relative to a reference stochastic process Léonard (2014). Variants of the SB problem have been applied across various domains, including single-cell RNA sequencing (scRNA-seq) analysis and generative modeling. These approaches encompass static methods Lavenant et al. (2024); Pariset et al. (2023); Ventre et al. (2023); Chizat et al. (2022); Shi et al. (2024); De Bortoli et al. (2021); Pooladian and Niles-Weed (2024); Liu et al. (2022); Gu et al. (2024), dynamic methods Koshizuka and Sato (2023); Neklyudov et al. (2023, 2024); Zhang et al. (2024); Bunne et al. (2023); Chen et al. (2022); Albergo et al. (2023); Wang et al. (2021); Jiao et al. (2024); Zhou et al. (2024); Liu et al. (2023); Zhou et al. (2024); Zhu et al. (2024); Maddu et al. (2024); Yeo et al. (2021); Jiang and Wan (2024), and flow-matching techniques Tong et al. (2024). However, these methods often fail to address unnormalized distributions resulting from cell growth and death. To address this problem, some methods have been developed to account for the unbalanced stochastic dynamics, for example, those based on branching SDE theory (e.g, gWOT) Chizat et al. (2022); Zhang et al. (2021); Ventre et al. (2023); Lavenant et al. (2024) and those based on the Feynman–Kac formula with forward-backward SDE theory Pariset et al. (2023). Among those, most methods often require prior knowledge of these processes, such as growth or death rates  Pariset et al. (2023); Lavenant et al. (2024); Chizat et al. (2022); Zhang et al. (2021) or depend on additional information like cell lineage data Ventre et al. (2023).
@@
Recently, regularized unbalanced optimal transport (RUOT), also known as unbalanced Schrödinger bridge Chen et al. (2022), has emerged as a promising approach for modeling stochastic unbalanced continuous dynamics Baradat and Lavenant (2021); Chen et al. (2022); Buze and Duong (2023); Janati et al. (2020). RUOT can be viewed as an unbalanced relaxation of the dynamic Schrödinger bridge formulation. For instance, ref. Baradat and Lavenant (2021) elucidates the connection between certain RUOT formulations and branching Schrödinger bridges. Meanwhile, a new deep learning framework (DeepRUOT) Zhang et al. (2025), has been developed to learn general RUOT and infer continuous unbalanced stochastic dynamics from sample data based on derived Fisher regularization forms without requiring prior knowledge.
@@
The primary objective now transforms to determine the dynamics described by Equations (2) and (3) from observed data, given unnormalized distributions at $T$ discrete time points where $\boldsymbol{x}_{i}\in\mathbb{R}^{d}\sim{\nu}_{i}$ for each fixed time point $i\in\{0,\ldots,T-1\}$. Note that solely satisfying Equation (3) does not admit a unique solution. Consequently, we need to ensure that the inferred dynamics also adhere to certain energy minimization principles. Building upon Equation (3), the problem can be categorized into four distinct scenarios:(1) ${g}(\boldsymbol{x},t)=0\text{ and }\boldsymbol{\sigma}(\boldsymbol{x},t)=0,$ (2) ${g}(\boldsymbol{x},t)=0\text{ and }\boldsymbol{\sigma}(\boldsymbol{x},t)\neq 0,$ (3) ${g}(\boldsymbol{x},t)\neq 0\text{ and }\linebreak\boldsymbol{\sigma}(\boldsymbol{x},t)=0,$ (4) ${g}(\boldsymbol{x},t)\neq 0\text{ and }\boldsymbol{\sigma}(\boldsymbol{x},t)\neq 0.$ Each of these cases is examined to systematically address the learning of the underlying dynamics.
@@
Dynamical Optimal Transport ($g=0,\sigma=0$) In this case, it means the dynamics do not account for unblancedness and stochastic, i.e., the cellular dynamics are governed by $\mathrm{d}{\boldsymbol{x}}_{t}=\boldsymbol{b}({\boldsymbol{x}}_{t},t)\mathrm{d}t$. Then we can use the dynamical optimal transport to model these dynamics, also known as the Benamou–Brenier formulation Benamou and Brenier (2000), which can be stated as follows:
@@
$\begin{gathered}\dfrac{1}{2}\mathcal{W}_{2}^{2}\left(\nu_{0},\nu_{1}\right)=\inf_{\left({p}(\boldsymbol{x},t),\boldsymbol{b}(\boldsymbol{x},t)\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\|\boldsymbol{b}(\boldsymbol{x},t)\|_{2}^{2}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,\\ \mbox{s.t. }\partial_{t}p+\nabla\cdot\left(\boldsymbol{b}(\boldsymbol{x},t)p\right)=0,\ p|_{t=0}=\nu_{0},p|_{t=1}=\nu_{1}.\end{gathered}$ (11)
@@
The inclusion of the factor $\frac{1}{2}$ on the left-hand side ensures that the Wasserstein distance has a more physically meaningful interpretation; for example, it represents the total action required to transport one distribution to another. In this formulation, probability distributions are connected through a deterministic transport equation. It has been demonstrated that this dynamic formulation is equivalent to static optimal transport problem (Equation (7)) when employing the cost function $c(\boldsymbol{x},\boldsymbol{y})=\|\boldsymbol{x}-\boldsymbol{y}\|_{2}^{2}$
Neural ODE Solver
@@
Numerous methodologies have been proposed to solve dynamical optimal transport or its variants numerically. The basic approach involves employing a neural network, denoted as $\boldsymbol{b}_{\theta}(\boldsymbol{x},t)$, to parameterize $\boldsymbol{b}(\boldsymbol{x},t)$, and subsequently utilizing the ordinary differential equations (ODEs) that govern particle trajectories. From Problem (11), it is evident that the optimization process must address two distinct loss components: the first pertains to the computation of an energy-related loss, while the second concerns the reconstruction error ($p(\boldsymbol{x},1)=\nu_{1}$).
@@
Regarding energy loss, the high-dimensional nature of the integral presents significant challenges due to the curse of dimensionality. To mitigate this issue, the integral is approximated using Monte Carlo integration and continuous normalizing flow (CNF) techniques Tong et al. (2020). The strategy involves performing integration along the particle trajectories dictated by the ODE, i.e.,
$\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\|\boldsymbol{b}(\boldsymbol{x},t)\|_{2}^{2}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t=\mathbb{E}_{\boldsymbol{x}_{0}\sim\nu_{0}}\int_{0}^{1}\frac{1}{2}\|\boldsymbol{b}(\boldsymbol{x}(t),t)\|_{2}^{2}\mathrm{d}t,$
@@
where $\boldsymbol{x}(t)$ satisfies the ODE $\frac{\mathrm{d}\boldsymbol{x}}{\mathrm{d}t}=\boldsymbol{b}(\boldsymbol{x},t),\boldsymbol{x}_{0}\sim\nu_{0}$. For the distribution reconstruction loss, the authors incorporate an additional penalizing constraint. By integrating these two loss components, there exists a sufficiently large $\lambda\geq 0$ such that Tong et al. (2020); Huguet et al. (2022)
@@
$\dfrac{1}{2}\mathcal{W}_{2}^{2}\left(\nu_{0},\nu_{1}\right)=\inf_{({p}(\boldsymbol{x},t),\boldsymbol{b}(\boldsymbol{x},t))}\mathbb{E}_{\boldsymbol{x}_{0}\sim\nu_{0}}\int_{0}^{1}\frac{1}{2}\|\boldsymbol{b}(\boldsymbol{x}(t),t)\|_{2}^{2}\mathrm{d}t+\lambda\mathcal{D}(p(\boldsymbol{x},1),\nu_{1}).$
@@
Based on this formulation, TrajectoryNet Tong et al. (2020) computes both the energy and the reconstruction errors by using neural ODE Chen et al. (2018) to parametrize the velocity $\boldsymbol{b}(\boldsymbol{x},t)$.
@@
Recently, conditional flow matching (CFM) presents another efficient dynamical OT solver especially in high dimensionality case Lipman et al. (2023); Tong et al. (2024); Liu et al. (2023); Albergo et al. (2023). Assume that the probability path ${p}(\boldsymbol{x},t)$ and the corresponding vector field $\boldsymbol{b}(\boldsymbol{x},t)$ generating it are known, and that ${p}(\boldsymbol{x},t)$ can be efficiently sampled. Under these conditions, a neural network $\boldsymbol{b}_{\theta}(\boldsymbol{x},t)$ can be trained to approximate $\boldsymbol{b}(\boldsymbol{x},t)$ by minimizing the flow matching (FM) objective:
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1),\boldsymbol{x}\sim{p}(\boldsymbol{x},t)}\|\boldsymbol{b}_{\theta}(\boldsymbol{x},t)-\boldsymbol{b}(\boldsymbol{x},t)\|_{2}^{2}.$
@@
However, this objective is computationally intractable when dealing with general source and target distributions. Consider the specific case of Gaussian marginal densities, defined as ${p}(\boldsymbol{x},t)=\mathcal{N}(\boldsymbol{x}|\boldsymbol{\mu}(t),\sigma(t)^{2}\boldsymbol{I})$. The corresponding unique vector field that generates this density from $\mathcal{N}(\boldsymbol{x}|\boldsymbol{\mu}(0),\sigma(0)^{2}\boldsymbol{I})$ is $\boldsymbol{b}(\boldsymbol{x},t)=\boldsymbol{\mu}^{\prime}(t)+\frac{\sigma(t)}{\sigma^{\prime}(t)}(\boldsymbol{x}-\boldsymbol{\mu}(t)),$ where $\mu^{\prime}(t)$ and $\sigma^{\prime}(t)$ means the time derivative Tong et al. (2024); Lipman et al. (2023). Now, assume the marginal probability trajectory ${p}(\boldsymbol{x},t)$ is a mixture of conditional probability paths $p(\boldsymbol{x},t|\boldsymbol{z})$. Specifically, this can be expressed as:
${p}(\boldsymbol{x},t)=\int p(\boldsymbol{x},t|\boldsymbol{z})q(\boldsymbol{z})\mathrm{d}\boldsymbol{z}.$
@@
If the $p(\boldsymbol{x},t|\boldsymbol{z})$ is generated by the vector field $\boldsymbol{b}(\boldsymbol{x},t|\boldsymbol{z})$ from $p(\boldsymbol{x},0|\boldsymbol{z})$, then ${p}(\boldsymbol{x},t)$ can be generated by $\boldsymbol{b}(\boldsymbol{x},t)$ defined as follows:
$\boldsymbol{b}(\boldsymbol{x},t):=\mathbb{E}_{q}(\boldsymbol{z})\frac{\boldsymbol{b}(\boldsymbol{x},t|\boldsymbol{z})p(\boldsymbol{x},0|\boldsymbol{z})}{{p}(\boldsymbol{x},t)}.$
@@
This is also intractable since ${p}(\boldsymbol{x},t)$ is difficult to compute. The key is to introduce the conditional flow matching objective:
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1),q(\boldsymbol{z}),p(\boldsymbol{x},t|\boldsymbol{z})}\|\boldsymbol{b}_{\theta}(\boldsymbol{x},t)-\boldsymbol{b}(\boldsymbol{x},t|\boldsymbol{z})\|_{2}^{2}.$
@@
One can prove that $\nabla_{\theta}\mathcal{L}_{\text{CFM}}=\nabla_{\theta}\mathcal{L}_{\text{FM}}$, so training with CFM is equivalent with FM. The CFM objective is very useful when the $\boldsymbol{b}(\boldsymbol{x},t)$ is intractable but the conditional $\boldsymbol{b}(\boldsymbol{x},t|\boldsymbol{z})$ is tractable. So to approximate the dynamical optimal transport (11) is to use CFM. Assume $q(\boldsymbol{z})=q(\boldsymbol{z}_{0},\boldsymbol{z}_{1})$, and set $q(\boldsymbol{z})$ to be the Wasserstein optimal transport map $\boldsymbol{\pi}$ between the source distribution $\nu_{0}$ and the target distribution $\nu_{1}$, i.e., $q(\boldsymbol{z})=\boldsymbol{\pi}(\boldsymbol{z}_{0},\boldsymbol{z}_{1})$, where $\boldsymbol{z}_{0}\sim\nu_{0}$, $\boldsymbol{z}_{1}\sim\nu_{1}$. Then one can construct the Gaussian flow between $\boldsymbol{z}_{0}$ and $\boldsymbol{z}_{1}$ with standard deviation $\sigma$,
$p(\boldsymbol{x},t|\boldsymbol{z})=\mathcal{N}(\boldsymbol{x}|t\boldsymbol{z}_{1}+(1-t)\boldsymbol{z}_{0}|\sigma^{2}),\boldsymbol{b}(\boldsymbol{x},t|\boldsymbol{z})=\boldsymbol{z}_{1}-\boldsymbol{z}_{0}.$